• What is the height of a projectile with initial velocity = "v" and angle = "theta" and gravity = "g"
1
Expert's answer
2021-07-12T12:16:10-0400

The height is given as follows (see https://courses.lumenlearning.com/boundless-physics/chapter/projectile-motion/):


"h = v\cdot t\cdot \sin \theta - \dfrac{gt^2}{2}"

where "t" is the time.